Your father offers you a choice of $135,000 in 10 years or $50,000 today. Use Appendix B as an approximate answer, but calculate your final answer using the formula and financial calculator methods.

a-1. If money is discounted at 9 percent, what is the present value of the $135,000? (Do not round intermediate calculations. Round your final answer to 2 decimal places.)

Present value


a-2. Which offer should you choose?
  

  • $50,000 today

  • $135,000 in 10 years

b-1. Now assume the offer is $135,000 in 7 years or $50,000 today. What is the present value of the $135,000 at 9 percent for 7 years? (Do not round intermediate calculations. Round your final answer to 2 decimal places.)

Present value


b-2. Now, which offer should you choose?
  

  • $135,000 in 7 years

  • $50,000 today
Answer #1

Option 1 $135000 in 10 years

Option 2 $50000 today,

discounting rate r = 9%

using PV = FV/(1+r)^t

a-1). PV of $135000 in 10 years = 135000/1.09^10 = $57025.46

a-2). As, PV of $135000 in 10 years is higher, this offer is selected.

When time is 7 years.

b-1).

a-1). PV of $135000 in 7 years = 135000/1.09^7 = $73849.62

a-2). As, PV of $135000 in 10 years is higher, this offer is selected.